# Functionally Constrained Algorithm Solves Convex Simple Bilevel Problems

### v2

@@ 7 Conclusion and future work
This paper provides a comprehensive study of convex simple bilevel problems. We show that find a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution for such problems is intractable for any zero-respecting first-order algorithm, thus justifying the notion of weak optimal solution considered by existing works. We then propose a novel framework FC-BiO for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Our proposed methods achieve the near-optmal rates of $\tilde{\mathcal{O}}\left(\max\{L_{f}^{2}/\epsilon_{f}^{2},L_{g}^{2}/\epsilon_{g}^{2}\}\right)$ and $\tilde{\mathcal{O}}\left(\max\{\sqrt{L_{f}/\epsilon_{f}},\sqrt{L_{g}/\epsilon_{g}}\}\right)$ for Lipschitz and smooth problems respectively.
@@ B.1 Determining $\hat{g}^{*}$ and the initial interval
To apply the two-step reformulation, we need to solve an approximate value of the lower-level problem $\hat{g}^{*}$ such that $g^{*}\leq\hat{g}^{*}\leq g^{*}+\frac{\epsilon}{2}$. Furthermore, to conduct the bisection method of FC-BiO, it is necessary to first determine an initial interval $[\ell,u]$ that contains the minimal root of $\psi^{*}(t)$, which is exactly $\hat{f}^{*}$ (Lemma 5.2). The following procedure is inspired by Wang et al. [27]: First, we apply optimal first-order methods for single-level optimization problems – subgradient method (SGM) [4, Section 3.1] for Lipschitz functions, and accelerated gradient method (AGM) [19, Section 2.2] for smooth functions – to the lower-level objective $g$ to obtain an approximate minimum point $\hat{\bf{x}}_{g}\in{\mathbb{R}}^{n}$ such that $g^{*}\leq g(\hat{\bf{x}}_{g})\leq g^{*}+\frac{\epsilon}{2}$. We set $\hat{g}^{*}=g(\hat{\bf{x}}_{g})$ and define the relaxed constraint function $\tilde{g}({\bf{x}})$ as in (4). We set $u=f(\hat{\bf{x}}_{g})$, then
@@
To derive a lower bound of the minimal root $\hat{f}^{*}$, we can also apply the optimal single-level minimization methods to the upper-level objective $f$ to find an approximate global minimum point $\hat{\bf{x}}_{f}\in{\mathbb{R}}^{n}$ such that $p^{*}\leq\hat{p}^{*}\triangleq f(\hat{\bf{x}}_{f})\leq p^{*}+\frac{\epsilon}{2}$, where $p^{*}$ is the minimum value of $f$ over the whole space ${\mathbb{R}}^{n}$ (i.e., $p^{*}\triangleq\min_{{\bf{x}}\in{\mathbb{R}}^{n}}f({\bf{x}})$). Then $\ell=\hat{p}^{*}-\frac{\epsilon}{2}$ is a valid lower bound of $\hat{f}^{*}$ since
@@ B.2 Warm-start strategy
For the initial point of the subroutine ${\mathcal{M}}$ (SGM in FC-BiOLip or generalized AGM in FC-BiOSm ), we exploit the warm-start strategy, which uses the last point in the previous round (denoted by $\bar{\bf{x}}$ in Algorithm 1) as the initial point of the current round. Intuitively, the subproblem (i.e. minimizing $\max\{f({\bf{x}})-t,g({\bf{x}})\}$) does not change significantly in each round since the parameter $t$ does not change too much as $\ell$ and $u$ are getting closer. Thus the approximate solution of the last round is supposed to be close to the optimal solution of the current round. Nevertheless, any initial point in ${\mathcal{B}}({\bf{x}}_{0},D)$ does not change the complexity upper bound. Such a strategy is widely used in two-level optimization methods [2, 16, 9].
@@ Appendix C Proofs for Section 5
Lemma C.1 (Nesterov [19, Theorem 2.1.7]).
@@
Lemma C.2 (Nesterov [19, Theorem 3.2.1]).
@@
$\bar{f}(t,{\bf{x}};{\bf{y}}_{k})=f({\bf{y}}_{k})-t-\frac{1}{2L}\|\nabla f({\bf{y}}_{k})\|^{2}+\frac{L}{2}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|^{2},$
$\bar{g}({\bf{x}};{\bf{y}}_{k})=\tilde{g}({\bf{y}}_{k})-\frac{1}{2L}\|\nabla g({\bf{y}}_{k})\|^{2}+\frac{L}{2}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|^{2}.$
@@
${\bf{x}}_{1}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{f}(t,{\bf{x}};{\bf{y}}_{k}),$
${\bf{x}}_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in{\mathcal{Z}}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{g}(t,{\bf{x}};{\bf{y}}_{k}).$
@@
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}\cap\mathcal{H}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|$
@@
$\min_{{\bf{x}}\in{\mathbb{R}}^{n}}f({\bf{x}}),\quad{\rm s.t.}\quad\tilde{g}^{\circ}({\bf{x}})=\frac{\epsilon_{f}}{\epsilon_{g}}\tilde{g}({\bf{x}})\leq 0.$ (15)
@@
Finally, we prove the claim at the end of Section 5 that our proposed algorithms are zero-respecting algorithms.
@@ Appendix D Finding absolute optimal solutions under additional assumptions
Proposition D.1 (Jiang et al. [12, Proposition 4.1]).
@@ E.1 Problem (11)
To be fair, all algorithms start from a random point ${\bf{x}}_{0}$ of unit length as the initial point. For our Algorithm 1, we take a slightly different implementation, that instead of setting the maximum number of iterations of the inner subroutine to be $T^{\prime}=T/N$, we preset $T^{\prime}=8000$. If current ${\bf{x}}_{k}$ already satisfies $\psi(t,{\bf{x}}_{k})\leq\frac{\epsilon}{2}$, then terminate the inner subroutine directly. We adopt the warm-start strategy as described in Appendix B.2. We set $L=0$ since $f({\bf{x}})$ is nonnegative. For FC-BiOLip, we set $\eta=3\times 10^{-4}$. For AGM-BiO, we set $\gamma=1$ as in [6, Theorem 4.1]. For PB-APG, we set $\gamma=10^{5}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{-3}$ and $\gamma_{0}=10^{-3}$. For CG-BiO, we set $\gamma_{k}=0.5/(k+1)$. For Bisec-BiO, we set the maximum number of iterations of the internal APG process to be $T^{\prime}=10000$. For FC-BiOsm, FC-BiOLip and Bisec-BiO, solving $\hat{g}^{*}$ takes $15000$ iterations; for CG-BiO, solving $\hat{g}^{*}$ takes $10000$ iterations. The results of such pretreatments are also plotted in Figure 1.

### v3

@@ 7 Conclusion and future work
This paper provides a comprehensive study of convex simple bilevel problems. We show that finding a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution for such problems is intractable for any zero-respecting first-order algorithm, thus justifying the notion of weak optimal solution considered by existing works. We then propose a novel method FC-BiO for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Our proposed method achieves the near-optmal rates of $\tilde{\mathcal{O}}\left(\max\{L_{f}^{2}/\epsilon_{f}^{2},L_{g}^{2}/\epsilon_{g}^{2}\}\right)$ and $\tilde{\mathcal{O}}\left(\max\{\sqrt{L_{f}/\epsilon_{f}},\sqrt{L_{g}/\epsilon_{g}}\}\right)$ for Lipschitz and smooth problems respectively. To the best of our knowledge, this is the first near-optimal algorithm that works under standard assumptions of smoothness or Lipschitz continuity for the objective functions.
@@ B.1 Determining $\hat{g}^{*}$ and the initial interval
To apply the two-step reformulation, we need to solve an approximate value of the lower-level problem $\hat{g}^{*}$ such that $g^{*}\leq\hat{g}^{*}\leq g^{*}+\frac{\epsilon}{2}$. Furthermore, to conduct the bisection method of FC-BiO, it is necessary to first determine an initial interval $[\ell,u]$ that contains the minimal root of $\psi^{*}(t)$, which is exactly $\hat{f}^{*}$ (Lemma 5.2). The following procedure is inspired by Wang et al. [28]: First, we apply optimal first-order methods for single-level optimization problems – subgradient method (SGM) [4, Section 3.1] for Lipschitz functions, and accelerated gradient method (AGM) [20, Section 2.2] for smooth functions – to the lower-level objective $g$ to obtain an approximate minimum point $\hat{\bf{x}}_{g}\in{\mathbb{R}}^{n}$ such that $g^{*}\leq g(\hat{\bf{x}}_{g})\leq g^{*}+\frac{\epsilon}{2}$. We set $\hat{g}^{*}=g(\hat{\bf{x}}_{g})$ and define the relaxed constraint function $\tilde{g}({\bf{x}})$ as in (4). We set $u=f(\hat{\bf{x}}_{g})$, then
@@
To derive a lower bound of the minimal root $\hat{f}^{*}$, we can also apply the optimal single-level minimization methods to the upper-level objective $f$ to find an approximate global minimum point $\hat{\bf{x}}_{f}\in{\mathcal{Z}}$ such that $p^{*}\leq\hat{p}^{*}\triangleq f(\hat{\bf{x}}_{f})\leq p^{*}+\frac{\epsilon}{2}$, where $p^{*}$ is the minimum value of $f$ over ${\mathcal{Z}}$ (i.e., $p^{*}\triangleq\min_{{\bf{x}}\in{\mathcal{Z}}}f({\bf{x}})$). Then $\ell=\hat{p}^{*}-\frac{\epsilon}{2}$ is a valid lower bound of $\hat{f}^{*}$ since
@@ B.2 Warm-start strategy
For the initial point of the subroutine ${\mathcal{M}}$ (SGM in FC-BiOLip or generalized AGM in FC-BiOSm ), we exploit the warm-start strategy, which uses the last point in the previous round (denoted by $\bar{\bf{x}}$ in Algorithm 1) as the initial point of the current round. Intuitively, the subproblem (i.e. minimizing $\max\{f({\bf{x}})-t,g({\bf{x}})\}$) does not change significantly in each round since the parameter $t$ does not change too much as $\ell$ and $u$ are getting closer. Thus the approximate solution of the last round is supposed to be close to the optimal solution of the current round. Nevertheless, any initial point in ${\mathcal{Z}}$ does not change the complexity upper bound. Such a strategy is widely used in two-level optimization methods [2, 17, 9].
@@ Appendix C Proofs for Section 5
Lemma C.1 (Nesterov [20, Theorem 2.1.7]).
@@
Lemma C.2 (Nesterov [20, Theorem 3.2.1]).
@@
$\bar{f}(t,{\bf{x}};{\bf{y}}_{k})=f({\bf{y}}_{k})-t-\frac{1}{2L}\|\nabla f({\bf{y}}_{k})\|^{2}+\frac{L}{2}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|_{2}^{2},$
$\bar{g}({\bf{x}};{\bf{y}}_{k})=\tilde{g}({\bf{y}}_{k})-\frac{1}{2L}\|\nabla g({\bf{y}}_{k})\|^{2}+\frac{L}{2}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|_{2}^{2}.$
@@
${\bf{x}}_{1}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{f}(t,{\bf{x}};{\bf{y}}_{k}),$
${\bf{x}}_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in{\mathcal{Z}}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{g}(t,{\bf{x}};{\bf{y}}_{k}).$
@@
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}\cap\mathcal{H}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|_{2}$
@@
$\min_{{\bf{x}}\in{\mathcal{Z}}}f({\bf{x}}),\quad{\rm s.t.}\quad\tilde{g}^{\circ}({\bf{x}})=\frac{\epsilon_{f}}{\epsilon_{g}}\tilde{g}({\bf{x}})\leq 0.$ (15)
@@
Finally, we prove the claim that our proposed algorithms are zero-respecting algorithms when the domain is a Eulidean ball centered at ${\bf{x}}_{0}$.
@@ Appendix D Finding absolute optimal solutions under additional assumptions
Proposition D.1 (Jiang et al. [13, Proposition 4.1]).
@@ E.1 Problem (11)
To be fair, all algorithms start from the same random point ${\bf{x}}_{0}$ of unit length as the initial point. For our Algorithm 1, we take a slightly different implementation, that instead of setting the maximum number of iterations of the inner subroutine to be $T^{\prime}=T/N$, we preset $T^{\prime}=8000$. If current ${\bf{x}}_{k}$ already satisfies $\psi(t,{\bf{x}}_{k})\leq\frac{\epsilon}{2}$, then terminate the inner subroutine directly. We adopt the warm-start strategy as described in Appendix B.2. We set $L=0$ since $f({\bf{x}})$ is nonnegative. For FC-BiOLip, we set $\eta=3\times 10^{-4}$. For AGM-BiO, we set $\gamma=1$ as in [6, Theorem 4.1]. For PB-APG, we set $\gamma=10^{5}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{-3}$ and $\gamma_{0}=10^{-3}$. For CG-BiO, we set $\gamma_{k}=0.5/(k+1)$. For Bisec-BiO, we set the maximum number of iterations of the internal APG process to be $T^{\prime}=10000$. For FC-BiOsm, FC-BiOLip and Bisec-BiO, solving $\hat{g}^{*}$ takes $15000$ iterations; for CG-BiO, solving $\hat{g}^{*}$ takes $10000$ iterations. The results of such pretreatments are also plotted in Figure 1.